# Amplitude equations of associative memory patterns in spatially distributed systems

### v2

@@ Abstract
Evolution equations of the amplitudes of heterogeneous states (associative memories) stored in the connectivity of distributed systems with non-local interactions are derived. The resulting system of coupled amplitude equations describes the spatio-temporal dynamics of memory recall. It is shown that these types of distributed systems perform pattern completion and selection, and that short-range connections afford spatio-temporal memory pattern dynamics taking the form of patterning fronts.
@@
Other processes, however, involve the formation of specific spatially heterogeneous states, with the spatial patterns stored in the heterogeneous interactions of the constituent parts: the recall of stored distributed memories. Memory storage and recall are commonly modelled using recurrently coupled neural associative memory networks [8, 9, 10, 11], of which modern iterations [12] drastically increase memory capacity [13, 14, 15, 16] and address biological validity [17, 18, 19].
In this letter amplitude equation derivation is generalised from spatially periodic modes to arbitrary spatially patterned states (memories) stored in the non-local connectivity of distributed systems. Hence, formalising a conjectured [20] link between the theory of pattern formation and dynamics, and associative memory networks.
The derivation naturally accommodates short-range interactions, which are common for physical and biological systems. The presented approach is therefore well-suited to study memory recall in systems with spatial constraints, such as biological neuronal networks [21, 22, 23]. The results demonstrate that associative memories can be stored using short-range interactions, and that this allows for novel spatio-temporal memory recall dynamics.
@@
$\gamma>\gamma_{k}=a_{1}\left[b_{1}\int_{\Omega}\int_{\Omega}J(\mathbf{r},\mathbf{r}^{\prime})\mu_{k}(\mathbf{r}^{\prime})d\mathbf{r}^{\prime}\mu_{k}(\mathbf{r})d\mathbf{r}\right]^{-1},$
@@
Amplitude equations.—The derivation of the amplitude equations is analogous to that for spatially periodic patterned states (see [3, 4]). For clarity of exposition it will be assumed that $u_{0}=0$, and that $f(u)$ and $g(u)$ are odd functions around $u_{0}$ so that their even-order derivatives vanish at the base-state $u_{0}$. Introducing the slow spatial $\tilde{\mathbf{r}}$ and temporal $\tilde{t}$ scales, and the perturbation expansion $U\sim U_{0}+\varepsilon U_{1}+\varepsilon^{2}U_{2}+\ldots$, with $U_{0}=u_{0}$ and $\varepsilon U_{1}=\sum_{k=1}^{M}A_{k}(\tilde{\mathbf{r}},\tilde{t})\mu_{k}(\mathbf{r})$, will give a system of $M$ coupled evolution equations for the pattern amplitudes $A_{k}$ as solvability conditions for the $O(\varepsilon^{3})$ terms.
@@
and at $O(\varepsilon^{3})$ the solvability criterion  [^3] Given that $\Phi$ is symmetric and, by assuming memory patterns with sufficient spatial variation on average, $J(\mathbf{r},\mathbf{r}^{\prime})$ will be symmetric around $\mathbf{r}^{\prime}=\mathbf{r}$, terms involving odd-order derivatives of $U_{1}$ in the expansion vanish under the integral.
@@
where $a_{3}=f^{\prime\prime\prime}(u_{0})/6$ and $b_{3}=g^{\prime\prime\prime}(u_{0})/6$. Finally, multiplying the $O(\varepsilon^{3})$ evolution equation of $U_{1}$ by $\mu_{k}$ and integrating over the domain gives a system of coupled amplitude equations (in original variables)
@@
Pattern dynamics.—The $M$ coupled equations, Eq. (3), describe the growth, interactions and (spatio-temporal) dynamics of the $M$ stored memory patterns in the full system. The coupled amplitude equations constitute a gradient system $\partial\mathbf{A}/\partial t=-\nabla V[\mathbf{A}]$, given the potential
@@
Bulk dynamics of the memories in the full system (grey markers in Fig. 1) are captured by the amplitude equations (black lines in Fig. 1), describing the pattern growth from low amplitude initial conditions, their transient dynamics, and the subsequent saturation of a single pattern to amplitude $A_{k}\to A_{k}^{\infty}=\varepsilon\sqrt{\alpha_{kk}/\beta_{kkkk}}$ together with the suppression of the other patterns $A_{m\neq k}\to 0$ due to the cubic interaction terms. For small enough $\varepsilon$ the coefficients of the linear terms $A_{k}$ and cubic interactions $A_{m}^{2}A_{k}$ in Eq. (3) will balance, so that in general the only stable equilibria will be single pattern states, as in the example of Fig. 1. Symmetric states, with multiple patterns having the same non-vanishing amplitude, are unstable equilibria.
@@
Spatial restriction of the connections ($\Phi$ vanishing for large arguments) leads to a finite diffusion rate of the memory patterns, and hence spatio-temporal pattern dynamics are expected in the full system. The equations governing the pattern amplitudes [Eq. (3)] are $M$ coupled parabolic partial differential equations, so travelling wave-front solutions, connecting $A_{k}=A_{k}^{\infty}$ behind the wave-front to $A_{k}=0$ infront of the wave-front, exist [27, 28, 29]. These solutions correspond to a patterned solution, $u(\mathbf{r},t)\propto\mu_{k}(\mathbf{r})$, invading an unpatterned spatially homogeneous domain (illustrated with the colour intensity in Fig. 2), in analogy to patterning fronts for spatially periodic states [30] and reflecting experimental evidence incidating travelling disturbances are involved in the recall and consolidation of memories brains [31]. The linear spreading velocity for a plane wave-front can be found by standard methods to be $c_{k}=\varepsilon\sqrt{2\alpha_{kk}D_{kk}}$. The theoretical front position, using $c_{k}$, is indicated with white dashed lines in Fig. 2 and agrees with the front in the full system.
Figure 2: Spatio-temporal profile of a pattern forming wave-front invading a spatially homogeneous domain. Colour intensity indicates the local pattern amplitude $u(r,t)\odot\mu_{k}(r)$, and the theoretical front position $r=r_{0}+\varepsilon\sqrt{2\alpha_{kk}D_{kk}}t$ is shown as white dashed lines.
@@
The ability of the derived amplitude equations to accurately capture the spatio-temporal dynamics of memory recall underscores a strong connection between pattern formation and associative memory. This invites the application of other techniques from pattern formation theory to study memory storage, recall, and selection in associative memory models and, given the natural handling of short-range connections, to the study memory in physical and biological systems.

### v3

@@ Abstract
Evolution equations are derived for the amplitudes of associative memories: heterogeneous states stored in the connectivity of distributed systems with non-local interactions. The resulting coupled amplitude equations describe the spatio-temporal dynamics of memory recall. They capture pattern completion and selection, and show that short-range connections can sustain spatio-temporal memory pattern dynamics in the form of propagating patterning fronts. The derived amplitude equations are of the same form as those describing classical pattern-forming instabilities, indicating a universality of the dynamics of memory recall and pattern formation in non-equilibrium systems.
@@
Other processes, however, involve the formation of specific spatially heterogeneous states, with the spatial patterns stored in the heterogeneous connectivity between the constituent parts: the recall of stored distributed memories. Memory storage and recall are commonly modelled using recurrently coupled neural associative memory networks [8, 9, 10, 11], of which modern iterations [12] drastically increase memory capacity [13, 14, 15, 16] and address biological validity [17, 18, 19].
In this letter amplitude equation derivation is generalised from spatially periodic modes to arbitrary spatially patterned states (memories) stored in the non-local connectivity of distributed systems. This suggests that memory recall and conventional pattern-forming instabilities belong to a common universality class, and hence formalises a conjectured [20, 21] link between associative memory networks, and the theory of pattern formation and dynamics.
The derivation naturally accommodates short-range interactions, making the presented approach well-suited to study memory recall in systems with spatial constraints, such as biological neuronal networks [22, 23, 24]. This aligns with previous results showing that associative memories can be stored using connections that decay with distance in some metric space [25, 26, 27, 28]. In addition to these results, the derived amplitude equations allow to study the emerging spatio-temporal memory recall dynamics.
@@
$\gamma>\gamma_{k}=a_{1}\left[b_{1}\left\langle\int_{\Omega}J(\mathbf{r},\mathbf{r}^{\prime})\mu_{k}(\mathbf{r}^{\prime})d\mathbf{r}^{\prime},\mu_{k}\right\rangle\right]^{-1},$
@@
Amplitude equations.—The derivation of the amplitude equations is analogous to that for spatially periodic patterned states (see [3, 4]). For clarity of exposition it will be assumed that $u_{0}=0$, and that $f(u)$ and $g(u)$ are odd functions around $u_{0}$ so that their even-order derivatives vanish at the base-state $u_{0}$. Introducing the slow spatial $\tilde{\mathbf{r}}$ and temporal $\tilde{t}$ scales, and the perturbation expansion $U\sim u_{0}+\varepsilon U_{1}+\varepsilon^{2}U_{2}+\ldots$, with $\varepsilon U_{1}=\sum_{k=1}^{M}A_{k}(\tilde{\mathbf{r}},\tilde{t})\mu_{k}(\mathbf{r})$, will give a system of $M$ coupled evolution equations for the pattern amplitudes $A_{k}$ as solvability conditions for the $O(\varepsilon^{3})$ terms as follows.
@@
and at $O(\varepsilon^{3})$ the solvability criterion  [^3] Given that $\Phi$ is symmetric and by assuming memory patterns with sufficient spatial variation on average so that $J(\mathbf{r},\mathbf{r}^{\prime})$ will be symmetric around $\mathbf{r}^{\prime}=\mathbf{r}$, terms involving odd-order derivatives of $U_{1}$ in the expansion vanish under the integral.
@@
where $a_{3}=f^{\prime\prime\prime}(u_{0})/6$ and $b_{3}=g^{\prime\prime\prime}(u_{0})/6$. Finally, multiplying the $O(\varepsilon^{3})$ evolution equation of $U_{1}$ by $\mu_{k}$ and integrating over the domain gives a system of coupled amplitude equations (in original variables):
@@
Pattern dynamics.—The $M$ coupled equations, Eq. (3), describe the growth, interactions and spatio-temporal dynamics of the $M$ stored memory patterns in the full system. The coupled amplitude equations constitute a gradient system $\partial\mathbf{A}/\partial t=-\nabla V[\mathbf{A}]$, given the potential
@@
Bulk dynamics of the memories in the full system (grey markers in Fig. 1) are captured by the amplitude equations (black lines in Fig. 1), describing the pattern growth from low amplitude initial conditions, their transient dynamics, and the subsequent saturation of a single pattern to amplitude $A_{k}\to A_{k}^{\infty}=\varepsilon\sqrt{\alpha_{kk}/\beta_{kkkk}}$ together with the suppression of the other patterns $A_{m\neq k}\to 0$ due to the cubic interaction terms. For small enough $\varepsilon$ the coefficients of the linear terms $A_{k}$ and cubic interactions $A_{m}^{2}A_{k}$ in Eq. (3) will balance, so that in general the only stable equilibria will be single pattern states, as in the example of Fig. 1. Symmetric states, with multiple patterns having the same non-vanishing amplitude in a single bulk domain, are unstable equilibria.
@@
Spatial restriction of the connections ($\Phi$ vanishing for large arguments) leads to a finite diffusion rate of the memory patterns, and hence spatio-temporal pattern dynamics are expected in the full system. The equations governing the pattern amplitudes [Eq. (3)] are $M$ coupled parabolic partial differential equations, so travelling wave-front solutions, connecting $A_{k}=A_{k}^{\infty}$ behind the wave-front to $A_{k}=0$ infront of the wave-front, exist [32, 33, 34]. These solutions correspond to a patterned solution, $u(\mathbf{r},t)\propto\mu_{k}(\mathbf{r})$, invading an unpatterned spatially homogeneous domain (illustrated with the colour intensity in Fig. 2), in analogy to patterning fronts for spatially periodic states [35] and reflecting experimental evidence incidating travelling disturbances are involved in the recall and consolidation of memories in brains [36]. The linear spreading velocity for a plane wave-front can be found by standard methods to be $c_{k}=\varepsilon\sqrt{2\alpha_{kk}D_{kk}}$. The theoretical front positions, using $c_{k}$, are indicated with white dashed lines in Fig. 2 and agree with the fronts in the full system.
Figure 2: Spatio-temporal profile of a pattern forming wave-front invading a spatially homogeneous domain. Colour intensity indicates the local pattern amplitude $\hat{a}_{k}(r,t)=2L\left[u(r,t)\odot\mu_{k}(r)\right]$. The theoretical front positions $r_{\pm}^{(k)}(t)=r_{\pm}^{(k)}(0)\pm\varepsilon\sqrt{2\alpha_{kk}D_{kk}}t$ are shown as white dashed lines.
Short-range connectivity also allows the contemporaneous retrieval of multiple different patterns in separated domains of the system [25, 26, 27, 28]: a state of multiple domains displaying different patterns separated by domain walls corresponds to a stable configuration of the system due to the normalisation of the memory patterns. In the case of memory patterns stored with differing strengths, spatio-temporal memory selection dynamics can emerge in the form of patterning fronts in which a strong memory invades the domain of a weaker memory.
@@
The ability of the derived amplitude equations to accurately capture the spatio-temporal dynamics of memory recall underscores a strong connection between pattern formation and associative memory. This invites the application of other insights from pattern formation theory to study memory storage, recall, and selection in associative memory models and, given the natural handling of short-range connections, to the study memory in physical and biological systems.